OpenStudy (drunblood):

Which fractions are equivalent to -3/8 ? Choose all answers that are correct. A. -6/16 B.12/32 C .-3/8 D. 9/-24

OpenStudy (drunblood):

need help please im new here but just thought yall would be of use so can i get help

jhonyy9 (jhonyy9):

than you multiplie top and bottom the -3/8 by the same number so will get fractions equivalent to given fraction of -3/8 for example (2/2)*(-3/8) = (-3*2)/(8*2) = -6/16 so result choice A is right sure hope you can continue it and this will help you good luck

OpenStudy (niuuu):

just a hint there are 2 more answers after the one jhonyy9 gave you